OpenStudy (emilyjones284):

A little help please...Ill give a medal (:

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

medal YUMYUM

OpenStudy (anonymous):

lol yummy^^

OpenStudy (anonymous):

YUMYUM !

OpenStudy (emilyjones284):

\[simplify: 2\sqrt[4]{2x}+6\sqrt[4]{2x}\]

OpenStudy (anonymous):

adding like terms

OpenStudy (anonymous):

\[(2 + 6) \sqrt[4]{2x}\] like what complete said just add up like terms

OpenStudy (emilyjones284):

ohhhhhh okay..got it thanks (:

OpenStudy (anonymous):

no problem :)